# Hybrid quantum lattice model: Polaritons, photons, and spin waves propagation

### v2

@@ III.3 Dispersive regime: Spin wave propagation
Spin-wave propagation under these conditions is illustrated in Fig. 4. Once released, the excitation propagates through the lattice via the TLS sublattice, forming a spin-wave [see Fig. 4(b)]. In this case, the average photon number in each resonator remains negligible across all sites [see Fig. 4(c)]. Figure 4(d) provides a global view of the excitation dynamics, showing the transport of the spin-wave through the hybrid lattice.
@@ IV Excitation Swap
Figure 5: (a) Schematic of the hybrid quantum lattice divided into two sections (left and right), each characterized by distinct parameter sets. (b) Schematic representation of the effective energy levels, effective coupling, and effective impedance-matching interface (green curve arrow), enabling an excitation swap of the type polariton $\rightarrow$ spin-wave can occur. The polariton transition frequencies and effective couplings between nearest-neighbour polaritons are indicated with purple arrows, while those corresponding to the spin-wave are shown in red. The orange arrow indicates the effective coupling between the activation qubit and the first polariton.
@@
Figure 6: Dynamic excitation swap, polariton $\rightarrow$ spin-wave: Average excitation number of the $j$th TLS and the average photon number as functions of time for the left section ($\ell$), shown in panels (a) and (b), and for the right section ($r$), shown in panels (c) and (d), respectively. (e) Average polariton number as a function of time and the site index $j$ (with $2\leq j\leq 31$). The parameters in the left section are $\omega_{\ell}=\omega^{\ell}_{0}=1$, $\Delta_{\ell}=\omega^{\ell}_{0}-\omega_{\ell}=0$, $v_{\ell}=0.05\,\omega_{\ell}$, $g_{\ell}=4v_{\ell}$, $\omega_{\rm{A}}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\lambda=\lambda_{\rm{C}}=-v_{\ell}/\sqrt{2}$. In the right section, the parameter are $\omega^{r}_{0}=\omega^{\ell}_{pol}$, $\omega_{r}=50\,\omega^{r}_{0}$, $\Delta_{r}=\omega^{r}_{0}-\omega_{r}$, $g_{r}=g_{\ell}$, $\theta_{1}=\arctan{(2g_{r}/\Delta_{r})}$, $v_{r}=v_{\ell}/2$. Additional parameters of the numerical simulation are $\tau=10^{-3}v_{\ell}^{-1}$, $n_{\rm{max}}=2$, and $L=31$.
@@
The hybrid quantum lattice allows controlled switching between excitation types through parameter tuning. As shown in the previous sections, depending on the regime (resonant or dispersive) and the specific conditions satisfied, the same injected excitation can propagate as a polariton, a photon, or a spin wave. This tunability enables a dynamic excitation swap, where the type of information carrier changes during propagation.
@@
To illustrate this mechanism, we divide the hybrid lattice into two sections, left and right, as shown in Fig. 5(a). All lattice parameters are labeled with $\ell$ and $r$ to indicate their respective sector. The excitation is injected into the left side via the activation qubit and propagates from left to right. The parameters are chosen such that the excitation propagates as one type on the left side and swaps to another type on the right side. The connection between the two regions is modeled by a coupling $\lambda_{\rm{C}}$, which acts analogously to an impedance-matching interface. Figure 5(b) shows a schematic representation of the effective energy levels, effective couplings, and the effective impedance-matching coupling. This configuration enables an excitation swap of the type polariton (left section) $\rightarrow$ spin-wave (right section).
@@
Note that $\lambda_{\rm{C}}$ is not dressed by the right side, since the selected excitation propagating in that region is a spin wave. For the excitation to be fully transmitted across the interface, the reflection on the left side must be minimized to near zero. This requires that the effective hopping of the polariton match the effective impedance-matching coupling, i.e., $\tilde{v}_{\ell}=\lambda_{\rm{eff}}$, where $\tilde{v}_{\ell}=v_{\ell}(\rho_{1-}^{\ell})^{2}$ [see Eq. 14]. The last expression implies the following
@@
Figure 6 illustrates the dynamic excitation swap from a polariton to a spin wave. This figure presents the average excitation number of the $j$th TLS, $\langle\hat{\sigma}^{+}_{j}\hat{\sigma}^{-}_{j}\rangle$ (upper row), and the average photon number, $\langle\hat{n}_{j}\rangle$ (lower row), as functions of time for each $j$-th resonator unit. Figures 6(a) and 6(b) show the dynamics of these two components of the average polariton number, $\langle\hat{\mathcal{N}}_{pol,j}\rangle$, on the left side of the lattice ($2\leq j\leq 15$). In this region, both components contribute equally to the propagation of the excitation pulse. As the pulse reaches the interface and enters the right section, both photon and TLS contributions in the left region diminish nearly to zero, indicating efficient transmission with no observable reflection. Figures 6(c) and 6(d) display the corresponding dynamics in the right section ($16\leq j\leq 31$). The excitation continues to propagate as a spin wave, as evidenced in Fig.6(c), while the photon number remains negligible, as shown in Fig.6(d). Figure 6(e) provides a complete view of the excitation dynamics across the entire system, showing the average polariton number $\langle\hat{\mathcal{N}}_{pol,j}\rangle$ as a function of time and site index. This global perspective confirms the efficient transfer of excitation from the polariton on the left to the spin wave on the right, with negligible reflection at the interface. Additionally, the propagation velocity remains unchanged across the two regions, as evidenced by the continuous linear slope of the excitation pulse.
@@
It is noteworthy that the excitation swap mechanism can also be realized in alternative hybrid lattice configurations. For instance, we consider a setup where the activation qubit is disconnected, and the system is initially prepared in a single polariton state localized at the center of the lattice ($j=(L+1)/2$ with $L$ odd). In this configuration, the lattice parameters are chosen such that the excitation propagates as a polariton to the left and as a spin wave to the right. The two pulses propagate in opposite directions, from the center toward the lattice boundaries, while conserving the total number of excitations, a single excitation. The impedance-matching condition between the two sections, $\lambda_{\rm{C}}$ is given by Eq. (22), and the coupling between TLSs in the right section, $v_{r}$, is determined by Eq. (24).
@@
Figure 7: Average polariton number, the average excitation number of the $j$th TLS and the average photon number as a function of time and the site index $j$ (with $1\leq j\leq 31$), shown in panels (a), (b), and (c), respectively. The parameters in the left section are $\omega_{\ell}=\omega^{\ell}_{0}=1$, $\Delta_{\ell}=\omega^{\ell}_{0}-\omega_{\ell}=0$, $v_{\ell}=0.01\,\omega_{\ell}$, $g_{\ell}=4v_{\ell}$, $\omega_{\rm{A}}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\lambda=\lambda_{\rm{C}}=-v_{\ell}/\sqrt{2}$. In the right section, the parameter are $\omega^{r}_{0}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\omega_{r}=50\,\omega^{r}_{0}$, $\Delta_{r}=\omega^{r}_{0}-\omega_{r}$, $g_{r}=g_{\ell}$, $\theta_{1}=\arctan{(2g_{r}/\Delta_{r})}$, $v_{r}=v_{\ell}/2$. Additional parameters of the numerical simulation are $\tau=10^{-3}v_{\ell}^{-1}$, $n_{\rm{max}}=2$, and $L=31$.
@@
Figure 7 shows the full dynamics of this process. Figure 7(a) shows the average polariton number as a function of time and site index, confirming the excitation’s symmetric splitting and propagation. Figure 7(b) displays the average excitation number of the $j$th TLS, highlighting the spin-wave character on the right half of the hybrid lattice. Figure 7(c) shows the average photon number, which remains negligible in the right section, further indicating the suppression of photonic components in the spin-wave propagation region.
V Conclusion
@@
In summary, our study demonstrates that a one-dimensional hybrid quantum lattice model, combining light-matter units with direct qubit-qubit interactions, enables controlled propagation and swapping of quantum excitations, namely, polaritons, spin waves, and photons. By adjusting system parameters to satisfy impedance and resonance conditions, we show how these excitations can be selectively guided or coherently swapped, with polaritons serving as the interface between photonic and spin-wave domains. For photon transport, we find that polaritonic states mediate single-excitation propagation, with negligible contribution from atomic excitations. This architecture provides a versatile platform for quantum links, where distant emitters can efficiently interact through polaritonic, spin-wave, or photonic networks, enabling robust quantum state transfer, hybrid information processing, and reconfigurable quantum networks. Future work should explore higher-excitation extensions, disorder effects in the hybrid lattice, and non-Markovian quantum links that communicate emitters.
@@ Acknowledgements.
M. A. acknowledges financial support from Postdoctoral FONDECYT Grant No. 3240443. We also thank the support from Dicyt USACH under grant $042331{\rm RH\textunderscore Ayudante}$, and Centro de Investigación Asociativa ANID CIA 250002.

### v3

@@ III.3 Dispersive regime: Atomic wave propagation
Atomic-wave propagation under these conditions is illustrated in Fig. 4. Once released, the excitation propagates through the lattice via the TLS sublattice, forming an atomic-wave [see Fig. 4(b)]. In this case, the average photon number in each resonator remains negligible across all sites [see Fig. 4(c)]. Figure 4(d) provides a global view of the excitation dynamics, showing the transport of the atomic-wave through the hybrid lattice.
@@ IV Excitation Swap
Figure 5: (a) Schematic of the hybrid quantum lattice divided into two sections (left and right), each characterized by distinct parameter sets. (b) Schematic representation of the effective energy levels, effective coupling, and effective impedance-matching interface (green curve arrow), enabling an excitation swap of the type polariton $\rightarrow$ atomic-wave can occur. The polariton transition frequencies and effective couplings between nearest-neighbour polaritons are indicated with purple arrows, while those corresponding to the atomic-wave are shown in red. The orange arrow indicates the effective coupling between the activation qubit and the first polariton.
@@
Figure 6: Dynamic excitation swap, polariton $\rightarrow$ atomic-wave: Average excitation number of the $j$th TLS and the average photon number as functions of time for the left section ($\ell$), shown in panels (a) and (b), and for the right section ($r$), shown in panels (c) and (d), respectively. (e) Average polariton number as a function of time and the site index $j$ (with $2\leq j\leq 31$). The parameters in the left section are $\omega_{\ell}=\omega^{\ell}_{0}=1$, $\Delta_{\ell}=\omega^{\ell}_{0}-\omega_{\ell}=0$, $v_{\ell}=0.05\,\omega_{\ell}$, $g_{\ell}=4v_{\ell}$, $\omega_{\rm{A}}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\lambda=\lambda_{\rm{C}}=-v_{\ell}/\sqrt{2}$. In the right section, the parameter are $\omega^{r}_{0}=\omega^{\ell}_{pol}$, $\omega_{r}=50\,\omega^{r}_{0}$, $\Delta_{r}=\omega^{r}_{0}-\omega_{r}$, $g_{r}=g_{\ell}$, $\theta_{1}=\arctan{(2g_{r}/\Delta_{r})}$, $v_{r}=v_{\ell}/2$. Additional parameters of the numerical simulation are $\tau=10^{-3}v_{\ell}^{-1}$, $n_{\rm{max}}=2$, $\chi_{\rm max}=50$, and $L=31$.
@@
Here, we describe the main contribution of our work, namely, the hybrid quantum lattice allows controlled swapping between excitation types through parameter tuning. As shown in the previous sections, depending on the regime (resonant or dispersive) and the specific conditions satisfied, the same injected excitation can propagate as a polariton, a photon, or an atomic wave. This tunability enables a dynamic excitation swap, where the type of information carrier changes during propagation.
@@
To illustrate this mechanism, we divide the hybrid lattice into two sections, left and right, as shown in Fig. 5(a). All lattice parameters are labeled with $\ell$ and $r$ to indicate their respective sector. The excitation is injected into the left side via the activation qubit and propagates from left to right. The parameters are chosen such that the excitation propagates as one type on the left side and swaps to another type on the right side. The connection between the two regions is modeled by a coupling $\lambda_{\rm{C}}$, which acts analogously to an impedance-matching interface. Figure 5(b) shows a schematic representation of the effective energy levels, effective couplings, and the effective impedance-matching coupling. This configuration enables an excitation swap of the type polariton (left section) $\rightarrow$ atomic-wave (right section).
@@
Note that $\lambda_{\rm{C}}$ is not dressed by the right side, since the selected excitation propagating in that region is an atomic wave. For the excitation to be fully transmitted across the interface, the reflection on the left side must be minimized to near zero. This requires that the effective hopping of the polariton match the effective impedance-matching coupling, i.e., $\tilde{v}_{\ell}=\lambda_{\rm{eff}}$, where $\tilde{v}_{\ell}=v_{\ell}(\rho_{1-}^{\ell})^{2}$ [see Eq. 14]. The last expression implies the following
@@
Besides, the resonance condition of the swap mechanism is given by
$\omega^{r}_{0}=\omega^{\ell}_{pol}\,$, (25)
indicating that the transition frequency of the lower polaritonic branch in the left lattice side must match the transition frequency of the right-side atoms [see Fig. 5(b)].
@@
Figure 6 illustrates the dynamic excitation swap from a polariton to an atomic wave. This figure presents the average excitation number of the $j$th TLS, $\langle\hat{\sigma}^{+}_{j}\hat{\sigma}^{-}_{j}\rangle$ (upper row), and the average photon number, $\langle\hat{n}_{j}\rangle$ (lower row), as functions of time for each $j$-th resonator unit. Figures 6(a) and 6(b) show the dynamics of these two components of the average polariton number, $\langle\hat{\mathcal{N}}_{pol,j}\rangle$, on the left side of the lattice ($2\leq j\leq 16$). In this region, both components contribute equally to the propagation of the excitation pulse. As the pulse reaches the interface and enters the right section, both photon and TLS contributions in the left region diminish nearly to zero, indicating efficient transmission with no observable reflection. Figures 6(c) and 6(d) display the corresponding dynamics in the right section ($17\leq j\leq 31$). The excitation continues to propagate as an atomic wave, as evidenced in Fig.6(c), while the photon number remains negligible, as shown in Fig.6(d). Figure 6(e) provides a complete view of the excitation dynamics across the entire system, showing the average polariton number $\langle\hat{\mathcal{N}}_{pol,j}\rangle$ as a function of time and site index. This global perspective confirms the efficient transfer of excitation from the polariton on the left to the atomic wave on the right, with negligible reflection at the interface. Additionally, the propagation velocity remains unchanged across the two regions, as evidenced by the continuous linear slope of the excitation pulse.
A sensitivity analysis of the swap mechanism is presented in Appendix D. There, we present results illustrating the swap dynamics when the matching Eq. (22) and resonance conditions Eq. (25) are intentionally detuned from their optimal values. Our results demonstrate that the swapping protocol is robust to moderate variations in the impedance-matching condition [see Fig. D1], while the resonance condition is more stringent: significant detuning leads to strong reflection and the suppression of excitation conversion [see Fig. D2].
As an additional diagnosis of the single excitation swap, we analyze entanglement via pairwise concurrences [63] between nearest-neighbor two-level systems, by tracing out photonic degrees of freedom. Given the reduced two-qubit density matrix $\rho_{j,j+1}$, the concurrence is defined as $C(\rho_{j,j+1})=\rm{max}(0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4})$ where $\{\lambda_{k}\}$ are the square roots of the eigenvalues of $R=\rho_{j,j+1}\,(\sigma_{y}\otimes\sigma_{y})\,\rho_{j,j+1}^{\ast}\,(\sigma_{y}\otimes\sigma_{y}),$ ordered as $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$, and $\sigma_{y}$ is the Pauli matrix. Figure 7 shows the pairwise concurrences as a function of time, where the upper panel shows concurrences of TLSs lying in the left part of the lattice. Concurrences exhibit similar patterns compared to local observables (c.f. Fig.6). This occurs due to the local interaction between the TLS and the photon. The highlighted concurrence (purple) corresponds to TLSs connecting both sides of the lattice. The lower panel shows the pairwise concurrences of TLSs lying in the right part of the lattice. Here we see an increase in the magnitude of concurrences since each TLS locally decouples from the photonic degrees of freedom, and entanglement is shared between nearest-neighbor TLSs only.
Figure 7: Pairwise concurrences: $C(\rho_{j,j+1})$ as a function of time for (a) $2\leq j\leq 16$ and (b) $17\leq j\leq 30$. The purple line shows the concurrence of the impedance-matching interface between the left and right lattice sections, $C(\rho_{16,17})$. The parameters are the same as Fig. 6.
@@
Interestingly, the excitation swap mechanism can also be realized in alternative hybrid lattice configurations. Let us consider a setup where the activation qubit is disconnected, and the system is initially prepared in a single polariton state localized at the center of the lattice ($j=(L+1)/2$ with $L$ odd). In this configuration, the lattice parameters are chosen such that the excitation propagates as a polariton to the left and as an atomic wave to the right. The two pulses propagate in opposite directions, from the center toward the lattice boundaries, while conserving the total number of excitations, a single excitation. The impedance-matching condition between the two sections, $\lambda_{\rm{C}}$ is given by Eq. (22), and the coupling between TLSs in the right section, $v_{r}$, is determined by Eq. (24).
@@
Figure 8: Average polariton number, the average excitation number of the $j$th TLS and the average photon number as a function of time and the site index $j$ (with $1\leq j\leq 31$), shown in panels (a), (b), and (c), respectively. The parameters in the left section are $\omega_{\ell}=\omega^{\ell}_{0}=1$, $\Delta_{\ell}=\omega^{\ell}_{0}-\omega_{\ell}=0$, $v_{\ell}=0.01\,\omega_{\ell}$, $g_{\ell}=4v_{\ell}$, $\omega_{\rm{A}}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\lambda=\lambda_{\rm{C}}=-v_{\ell}/\sqrt{2}$. In the right section, the parameter are $\omega^{r}_{0}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\omega_{r}=50\,\omega^{r}_{0}$, $\Delta_{r}=\omega^{r}_{0}-\omega_{r}$, $g_{r}=g_{\ell}$, $\theta_{1}=\arctan{(2g_{r}/\Delta_{r})}$, $v_{r}=v_{\ell}/2$. Additional parameters of the numerical simulation are $\tau=10^{-3}v_{\ell}^{-1}$, $n_{\rm{max}}=2$, $\chi_{\rm max}=4$, and $L=31$.
@@
Figure 8(a) shows the average polariton number as a function of time and site index, confirming the excitation’s symmetric splitting and propagation. Figure 8(b) displays the average excitation number of the $j$th TLS, highlighting the atomic-wave character on the right half of the hybrid lattice. Figure 8(c) shows the average photon number, which remains negligible in the right section, further indicating the suppression of photonic components in the atomic-wave propagation region.
V Experimental Feasibility in Circuit-QED Platforms
To assess the practical relevance of the proposed excitation-swapping mechanism, we now discuss its implementation in state-of-the-art superconducting circuit-QED platforms. In such architectures, tight-binding lattice sites can be realized using transmon qubits, while the effective tunneling between sites arises from capacitive coupling between neighboring qubits [37, 66, 67, 68, 76, 77, 78, 79]. The on-site energies can be tuned with site-resolved precision via local flux-bias lines, enabling full control over the lattice energy landscape. In addition, each transmon can be coupled to an off-resonant coplanar waveguide resonator, which mediates the light–matter interaction required to realize the modified Jaynes-Cummings-Hubbard model considered here. These capabilities suggest that the impedance-matching and resonance conditions introduced in our theoretical framework can be implemented experimentally.
In particular in Ref. [79], the authors report the experimentally achievable parameter ranges such as: transmon tuning range frequency $\omega_{0}/2\pi\in[3,6]\,\mathrm{GHz}$, capacitive coupling between nearest-neighbor transmon sites sets a tunneling rate $J/2\pi\approx 9\,\mathrm{MHz}$, qubit-resonator dispersive couplings range $\chi_{qR}/2\pi\in[0.5,1.6]\,\mathrm{MHz}$, qubit relaxation time $T_{1}\approx 45\,\mathrm{\mu s}$, and resonator linewidths $\kappa_{R}/2\pi\approx 100\,\mathrm{kHz}$. Using parameters within these experimentally reported ranges, the results shown in Fig. 6 can be reproduced, and the excitation-swapping mechanism remains robust under realistic conditions.
In our numerical simulations, the excitation swap occurs at a characteristic dimensionless time $t_{max}=35$. Converting to physical units using the experimentally relevant coupling $J/2\pi\approx 9\,\mathrm{MHz}$, this corresponds to a timescale $T_{max}=t_{max}/J=35/2\pi\times 9\,\mathrm{MHz}\approx 0.6\,\mathrm{\mu s}$. This should be compared with experimentally reported qubit relaxation times of $T_{1}\approx 45\,\mathrm{\mu s}$ in circuit-QED platform [79]. Since $T_{max}\ll T_{1}$, the excitation-swapping dynamics takes place well within the coherence time of the system. This separation of timescales indicates that the proposed many-body dynamical swapping effect is, in principle, observable before decoherence significantly affects the system.
In addition, using resonator linewidths $\kappa_{R}/2\pi\approx 100\,\mathrm{kHz}$ reported in Ref. [79], the photon lifetime in the resonators is $\tau_{ph}\approx 1.6\,\mathrm{\mu s}$. Since the characteristic timescale of the excitation swap in our simulations is $T_{max}=0.6\,\mathrm{\mu s}$, we have $T_{max}<\tau_{ph}$, indicating that the photonic component of the excitation remains coherent during the swapping process. Together with the condition $T_{max}\ll T_{1}$, this supports the feasibility of observing the proposed mechanism within current circuit-QED platforms.
Figure 9: Excitation swap using experimental parameters: Average polariton number as a function of time and site index $j$ (with $2\leq j\leq 31$). The parameters in the left section are $\omega_{\ell}=\omega^{\ell}_{0}=1$, $\Delta_{\ell}=\omega^{\ell}_{0}-\omega_{\ell}=0$, $v_{\ell}=0.003$, $g_{\ell}=4v_{\ell}$, $\omega_{\rm{A}}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\lambda=\lambda_{\rm{C}}=-v_{\ell}/\sqrt{2}$. In the right section, the parameters are $\omega^{r}_{0}=\omega^{\ell}_{pol}$, $\Delta_{r}=0.27$, $\omega_{r}=0.718$, $g_{r}=g_{\ell}$, $\theta_{1}=\arctan{(2g_{r}/\Delta_{r})}$, $v_{r}=v_{\ell}/2$. Additional parameters of the numerical simulation are $\tau=10^{-3}v_{\ell}^{-1}$, $n_{\rm{max}}=2$, $\chi_{\rm max}=50$, and $L=31$.
To further support the experimental feasibility, we report the numerical simulation using parameters within experimentally reported ranges. Figure 9 shows the average polariton number $\langle\hat{\mathcal{N}}_{pol,j}\rangle$ as a function of time and site index. The simulations are carried out using $\omega_{0}/2\pi=3\,\mathrm{GHz}$, $J/2\pi=9\,\mathrm{MHz}$, and $\chi_{qR}/2\pi=1.6\,\mathrm{MHz}$. All parameters in Fig. 9 are normalized with respect to $\omega_{0}/2\pi$. These results show that the excitation-swapping mechanism remains robust under realistic conditions, thereby reinforcing its potential applicability in circuit-QED platforms.
VI Conclusion
@@
In summary, our study demonstrates that a one-dimensional hybrid quantum lattice model, combining light-matter units with direct qubit-qubit interactions, enables controlled propagation and swapping of quantum excitations, namely, polaritons, atomic waves, and photons. By adjusting system parameters to satisfy impedance and resonance conditions, we show how these excitations can be selectively guided or coherently swapped, with polaritons serving as the interface between photonic and atomic-wave domains. For photon transport, we find that polaritonic states mediate single-excitation propagation, with negligible contribution from atomic excitations. This architecture provides a versatile platform for quantum links, where distant emitters can efficiently interact through polaritonic, atomic-wave, or photonic networks, enabling hybrid information processing and reconfigurable quantum networks. Future work should explore higher-excitation extensions, disorder effects in the hybrid lattice, and non-Markovian quantum links that communicate emitters.
@@ Acknowledgements.
M. A. acknowledges financial support from ANID Postdoctoral FONDECYT Grant No. 3240443. We also thank the support from Dicyt USACH under grant $042331{\rm RH\textunderscore Ayudante}$, Vicerrectoría de Investigación, Innovación y Creación, the Center for Nanoscience and Nanotechnology, CEDENNA, Project CIA250002, and Proyecto de Exploración Grant No. 13250014.